Numbe5

0 5. Find potential energy of a system of three poics charges shones in the tre 4 em -10 nC +15 AC
Answer #1

5)

Potential energy of +15uC due to -10uC ,

V1 = k*Q1*Q2/r

So, V1 = 9*10^9*(+15*10^-6)*(-10*10^-6)/0.04

Similarly Potential energy of +15uC due to +5uC

V2 = 9*10^9*(15*10^-6)*(5*10^-6)/(0.05) <------- distance between +15 uC and +5uC = 5 cm = 0.05 m

Now,

Potential energy of -10uC due to +5uC

V3 = 9*10^9*(-10*10^-6)*(5*10^-6)/(0.03)

Now, these are the only interactions between the charges possible for 3 charges,

So, Vnet = V1 +V2 +V3

So, Vnet = 9*10^9*(15*10^-6)*(-10*10^-6)/0.04 + 9*10^9*(15*10^-6)*(5*10^-6)/(0.05) + 9*10^9*(-10*10^-6)*(5*10^-6)/(0.03)

So, Vnet = -35.25 V